\begin{thebibliography}{100}

\bibitem{sornette1987}
{Gilabert, A.}, {Vanneste, C.}, {Sornette, D.}, and {Guyon, E.}, ``The random
  fuse network as a model of rupture in a disordered medium,'' {\em J. Phys.
  France}, vol.~48, no.~5, 1987.

\bibitem{0022-3727-42-21-214014}
D.~Bonamy, ``Intermittency and roughening in the failure of brittle
  heterogeneous materials,'' {\em Journal of Physics D Applied Physics},
  vol.~42, no.~21, p.~214014, 2009.

\bibitem{alava09}
M.~J. Alava, P.~K. V.~V. Nukala, and S.~Zapperi, ``Size effects in statistical
  fracture,'' {\em Journal of Physics D Applied Physics}, vol.~42, no.~21,
  p.~214012, 2009.

\bibitem{Alava2002}
L.~I. Salminen, A.~I. Tolvanen, and M.~J. Alava, ``Acoustic emission from paper
  fracture,'' {\em Phys. Rev. Lett.}, vol.~89, p.~185503, 2002.

\bibitem{alava2006}
M.~J. Alava, P.~K. V.~V. Nukala, and S.~Zapperi, ``Statistical models of
  fracture,'' {\em Advances in Physics}, vol.~55, no.~3, 2006.

\bibitem{alava2006morphology}
M.~J. Alava, P.~K. Nukala, and S.~Zapperi, ``Morphology of two-dimensional
  fracture surfaces,'' {\em Journal of Statistical Mechanics Theory and
  Experiment}, vol.~2006, no.~10, p.~L10002, 2006.

\bibitem{alava2008fracture}
M.~J. Alava, P.~K. Nukala, and S.~Zapperi, ``Fracture size effects from
  disordered lattice models,'' {\em International journal of fracture},
  vol.~154, no.~1-2, pp.~51--59, 2008.

\bibitem{alava2008role}
M.~Alava, P.~Nukala, and S.~Zapperi, ``Role of disorder in the size scaling of
  material strength,'' {\em Physical review letters}, vol.~100, no.~5,
  p.~55502, 2008.

\bibitem{Alava2010}
L.~Laurson, M.-C. Miguel, and M.~J. Alava, ``Dynamical correlations near
  dislocation jamming,'' {\em Phys. Rev. Lett.}, vol.~105, p.~015501, 2010.

\bibitem{Anup2015168}
S.~Anup, ``Influence of initial flaws on the mechanical properties of nacre,''
  {\em Journal of the Mechanical Behavior of Biomedical Materials}, vol.~46,
  no.~0, pp.~168 -- 175, 2015.

\bibitem{arcangelis85}
L.~de~Arcangelis, S.~Redner, and H.~J. Hermann, ``A random fuse model for
  breaking processes,'' {\em J.Physique Lett.}, vol.~46, pp.~L--585, 1985.

\bibitem{bardley1994}
R.~M. Bradley and K.~Wu, ``Dynamic fuse model for electromigration failure of
  polycrystalline metal films,'' {\em Phys. Rev. E}, vol.~50, pp.~R631--R634,
  1994.

\bibitem{bastiaansen97}
P.~J.~M. Bastiaansen and H.~J.~F. Knops, ``Correlated percolation and the
  correlated resistor network,'' {\em J. Phys. A Math. Gen.}, vol.~30, p.~1791,
  1997.

\bibitem{beale1988}
P.~D. Beale and P.~M. Duxbury, ``Theory of dielectric breakdown in metal-loaded
  dielectrics,'' {\em Phys. Rev. B}, vol.~37, no.~6, pp.~2785--2791, 1988.

\bibitem{bergman1981}
D.~J. Bergman and Y.~Kantor, ``Series expansion for the macroscopic
  conductivity of a random resistor network,'' {\em Journal of Physics C Solid
  State Physics}, vol.~14, p.~3365, 1981.

\bibitem{blackman76}
J.~A. Blackman, ``A theory of conductivity in disordered resistor networks,''
  {\em J. Phys. C Solid State Phys.}, vol.~9, p.~2049, 1976.

\bibitem{bour2002statistical}
O.~Bour, P.~Davy, C.~Darcel, and N.~Odling, ``A statistical scaling model for
  fracture network geometry, with validation on a multiscale mapping of a joint
  network hornelen basin, norway,'' {\em Journal of Geophysical Research},
  vol.~107, no.~B6, p.~2113, 2002.

\bibitem{BrownHCMLSC04}
K.~S. Brown, C.~C. Hill, G.~A. Calero, C.~R. Myers, K.~H. Lee, J.~P. Sethna,
  and R.~A. Cerione, ``The statistical mechanics of complex signaling networks
  nerve growth factor signaling,'' {\em Physical Biology}, vol.~1,
  pp.~184--195, 2004.

\bibitem{ChenPSZD11}
Y.~Chen, S.~Papanikolaou, J.~P. Sethna, S.~Zapperi, and G.~Durin, ``Avalanche
  spatial structure and multivariable scaling functions sizes, heights, widths,
  and views through windows,'' {\em Physical Review E}, vol.~84, p.~061103,
  2011.

\bibitem{shekhawat2012ProcI}
A.~{Shekhawat}, ``{Fracture Statistics: Universality vs. Nucleation},'' in {\em
  APS Meeting Abstracts}, p.~43004, Feb. 2012.

\bibitem{colina1993model}
H.~Colina, L.~de~Arcangelis, and S.~Roux, ``Model for surface cracking,'' {\em
  Physical Review B}, vol.~48, no.~6, pp.~3666--3676, 1993.

\bibitem{Csikor2007}
F.~F. Csikor, C.~Motz, D.~Weygand, M.~Zaiser, and S.~Zapperi, ``Dislocation
  avalanches, strain bursts, and the problem of plastic forming at the
  micrometer scale,'' {\em Science}, vol.~318, no.~5848, pp.~251--254, 2007.

\bibitem{curtin1990brittle}
W.~Curtin and H.~Scher, ``Brittle fracture in disordered materials a spring
  network model,'' {\em Journal of Materials Research}, vol.~5, no.~3,
  pp.~535--553, 1990.

\bibitem{davy1995localization}
P.~Davy, A.~Hansen, E.~Bonnet, and S.~Zhang, ``Localization and fault growth in
  layered brittle-ductile systems implications for deformations of the
  continental lithosphere,'' {\em Journal of Geophysical Research}, vol.~100,
  no.~B4, pp.~6281--6294, 1995.

\bibitem{de1989scaling}
L.~De~Arcangelis and H.~Herrmann, ``Scaling and multiscaling laws in random
  fuse networks,'' {\em Physical Review B}, vol.~39, no.~4, pp.~2678--2684,
  1989.

\bibitem{shekhawat2012ProcII}
A.~{Shekhawat}, S.~{Zapperi}, and J.~{Sethna}, ``{Fracture In Disordered Media:
  Nucleated, Critical or Percolative?},'' in {\em APS Meeting Abstracts},
  p.~26007, Feb. 2012.

\bibitem{de1989scalingI}
L.~De~Arcangelis, A.~Hansen, H.~Herrmann, and S.~Roux, ``Scaling laws in
  fracture,'' {\em Physical Review B}, vol.~40, no.~1, pp.~877--880, 1989.

\bibitem{delaplace1996progressive}
A.~Delaplace, G.~Pijaudier-Cabot, and S.~Roux, ``Progressive damage in discrete
  models and consequences on continuum modelling,'' {\em Journal of the
  Mechanics and Physics of Solids}, vol.~44, no.~1, pp.~99--136, 1996.

\bibitem{duxbury1986}
P.~M. Duxbury, P.~D. Beale, and P.~L. Leath, ``Size effect of electrical
  breakdown in quenched random media,'' {\em Phys. Rev. Lett.}, vol.~57, no.~8,
  1986.

\bibitem{duxbury1987}
P.~M. Duxbury, P.~L. Leath, and P.~D. Beale, ``Breakdown properties of quenched
  random systems the random-fuse network,'' {\em Phys. Rev. B}, vol.~36, no.~1,
  pp.~367--380, 1987.

\bibitem{duxbury1987failure}
P.~Duxbury and P.~Leath, ``The failure distribution in percolation models of
  breakdown,'' {\em Journal of Physics A Mathematical and General}, vol.~20,
  p.~L411, 1987.

\bibitem{duxbury1994}
P.~M. Duxbury and P.~L. Leath, ``Exactly solvable models of material
  breakdown,'' {\em Phys. Rev. B}, vol.~49, pp.~12676--12687, 1994.

\bibitem{duxbury1994failure}
P.~Duxbury and P.~Leath, ``Failure probability and average strength of
  disordered systems,'' {\em Physical review letters}, vol.~72, no.~17,
  pp.~2805--2808, 1994.

\bibitem{shekhawat2012ProcIII}
A.~Shekhawat, C.~Manzato, P.~Nukala, M.~Alava, S.~Zapperi, and J.~Sethna,
  ``Statistics of fracture: Weibull, gumbel and other questions,'' in {\em AIP
  Conference Proceedings}, American Institute of Physics, Ste. 1 NO 1 Melville
  NY 11747-4502 United States, 2012.

\bibitem{duxbury1994size}
P.~Duxbury, S.~Kim, and P.~Leath, ``Size effect and statistics of fracture in
  random materials,'' {\em Materials Science and Engineering A}, vol.~176,
  no.~1-2, pp.~25--31, 1994.

\bibitem{hansen1991}
S.~Roux, A.~Hansen, E.~L. Hinrichsen, and D.~Sornette, ``Fuse model on a
  randomly diluted hierarchical lattice,'' {\em Journal of Physics A
  Mathematical and General}, vol.~24, no.~7, p.~1625, 1991.

\bibitem{hansen1991roughness}
A.~Hansen, E.~Hinrichsen, and S.~Roux, ``Roughness of crack interfaces,'' {\em
  Physical review letters}, vol.~66, no.~19, pp.~2476--2479, 1991.

\bibitem{hansen2003}
A.~Hansen and J.~Schmittbuhl, ``Origin of the universal roughness exponent of
  brittle fracture surfacesstress-weighted percolation in the damage zone,''
  {\em Phys. Rev. Lett.}, vol.~90, p.~045504, 2003.

\bibitem{hansen2005}
A.~Hansen, ``Physics and fracture,'' {\em Computing in Science Engineering},
  vol.~7, no.~5, 2005.

\bibitem{hansen2006}
R.~Toussaint and A.~Hansen, ``Mean-field theory of localization in a fuse
  model,'' {\em Phys. Rev. E}, vol.~73, p.~046103, 2006.

\bibitem{hansen2012}
A.~A. Moreira, C.~L.~N. Oliveira, A.~Hansen, N.~A.~M. Araujo, H.~J. Herrmann,
  and J.~S. Andrade, ``Fracturing highly disordered materials,'' {\em Phys.
  Rev. Lett.}, vol.~109, p.~255701, 2012.

\bibitem{hemmer1992}
P.~C. Hemmer and A.~Hansen, ``The distribution of simultaneous fiber failures
  in fiber bundles,'' {\em Journal of applied mechanics}, vol.~59, p.~909,
  1992.

\bibitem{hemmer1997}
M.~Kloster, A.~Hansen, and P.~C. Hemmer, ``Burst avalanches in solvable models
  of fibrous materials,'' {\em Phys. Rev. E}, vol.~56, pp.~2615--2625, 1997.

\bibitem{hidalgo2001bursts}
R.~Hidalgo, F.~Kun, and H.~Herrmann, ``Bursts in a fiber bundle model with
  continuous damage,'' {\em Physical Review E}, vol.~64, no.~6, p.~66122, 2001.

\bibitem{shekhawat2014Imp}
A.~Shekhawat, ``Improving extreme value statistics,'' {\em Phys. Rev. E},
  vol.~90, p.~052148, Nov 2014.

\bibitem{Hope20094593}
S.~M. Hope and A.~Hansen, ``Burst distribution in noisy fiber bundles and fuse
  models,'' {\em Physica A Statistical Mechanics and its Applications},
  vol.~388, no.~21, pp.~4593 -- 4599, 2009.

\bibitem{HughChrzLiu98}
D.~A. Hughes, D.~C. Chrzan, Q.~Liu, and N.~Hansen, ``Scaling of misorientation
  angle distributions,'' {\em Phys. Rev. Lett.}, vol.~81, pp.~4664--4667, 1998.

\bibitem{HughesActaMater1997}
D.~A. Hughes, Q.~Liu, D.~C. Chrzan, and N.~Hansen, ``Scaling of microstructural
  parameters misorientations of deformation induced boundaries,'' {\em Acta
  materialia}, vol.~45, no.~1, pp.~105--112, 1997.

\bibitem{HughesPRL2001}
D.~A. Hughes and N.~Hansen, ``Graded nanostructures produced by sliding and
  exhibiting universal behavior,'' {\em Phys. Rev. Lett.}, vol.~87, no.~13,
  p.~135503, 2001.

\bibitem{HughesScrMater2003}
D.~A. Hughes, N.~Hansen, and D.~J. Bammann, ``Geometrically necessary
  boundaries, incidental dislocation boundaries and geometrically necessary
  dislocations,'' {\em Scripta Materialia}, vol.~48, no.~2, pp.~147 -- 153,
  2003.

\bibitem{JohansenApJ2006}
A.~Johansen, T.~Henning, and H.~Klahr, ``Dust sedimentation and self-sustained
  kelvin-helmholtz turbulence in protoplanetary disk midplanes,'' {\em The
  Astrophysical Journal}, vol.~643, p.~1219, 2006.

\bibitem{Kahng1988}
B.~Kahng, G.~G. Batrouni, S.~Redner, L.~de~Arcangelis, and H.~J. Herrmann,
  ``Electrical breakdown in a fuse network with random, continuously
  distributed breaking strengths,'' {\em Phys. Rev. B}, vol.~37,
  pp.~7625--7637, 1988.

\bibitem{korteoja1998}
M.~Korteoja, L.~I. Salminen, K.~J. Niskanen, and M.~J. Alava, ``Strength
  distribution in paper,'' {\em Materials Science and Engineering A}, vol.~248,
  no.~1-2, pp.~173 -- 180, 1998.

\bibitem{KuhlmannWilsdorfScrMetalMater1991}
D.~Kuhlmann-Wilsdorf and N.~Hansen, ``Geometrically necessary, incidental and
  subgrain boundaries,'' {\em Scripta Metallurgica et Materialia}, vol.~25,
  no.~7, pp.~1557 -- 1562, 1991.

\bibitem{kun2000damage}
F.~Kun, S.~Zapperi, and H.~Herrmann, ``Damage in fiber bundle models,'' {\em
  The European Physical Journal B-Condensed Matter and Complex Systems},
  vol.~17, no.~2, pp.~269--279, 2000.

\bibitem{LaPortaZS12}
C.~A. M.~L. Porta, S.~Zapperi, and J.~P. Sethna, ``Senescent cells in growing
  tumors population dynamics and cancer stem cells,'' {\em PLoS Computational
  Biology}, vol.~8, p.~e1002316, 2012.

\bibitem{LaursonA2006}
L.~Laurson and M.~J. Alava, ``1f noise and avalanche scaling in plastic
  deformation,'' {\em Phys. Rev. E}, vol.~74, p.~066106, 2006.

\bibitem{LuH2009}
K.~Lu and N.~Hansen, ``Structural refinement and deformation mechanisms in
  nanostructured metals,'' {\em Scripta Materialia}, vol.~60, no.~12, pp.~1033
  -- 1038, 2009.

\bibitem{LuoZHL2012}
Z.~Luo, H.~Zhang, N.~Hansen, and K.~Lu, ``Quantification of the microstructures
  of high purity nickel subjected to dynamic plastic deformation,'' {\em Acta
  Materialia}, vol.~60, no.~3, pp.~1322 -- 1333, 2012.

\bibitem{MagniDZS09}
A.~Magni, G.~Durin, S.~Zapperi, and J.~P. Sethna, ``Visualization of avalanches
  in magnetic thin films temporal processing,'' {\em Journal of Statistical
  Mechanics Theory and Experiment}, vol.~2009, no.~01, p.~P01020, 2009.

\bibitem{MagniDZSnn}
A.~Magni, D.~G., Z.~S., and J.~P. Sethna, ``Avalanches through windows
  multiscale visualization in magnetic thin films,'' {\em
  url{httpwww.lassp.cornell.edusethnapubPDFIntermag08Win.pdf} submitted}.

\bibitem{ManzatoSNASZ12}
C.~Manzato, A.~Shekhawat, P.~K. V.~V. Nukala, M.~J. Alava, J.~P. Sethna, and
  S.~Zapperi, ``Fracture strength of disordered media universality,
  interactions and tail asymptotics,'' {\em Physical Review Letters}, vol.~108,
  p.~065504, 2012.

\bibitem{shekhawat2013ProcI}
A.~{Shekhawat}, S.~{Zapperi}, and J.~{Sethna}, ``{Brittle Fracture In
  Disordered Media: A Unified Theory},'' in {\em APS Meeting Abstracts},
  p.~28009, Mar. 2013.

\bibitem{Miguel2001}
M.~C. Miguel, A.~Vespignani, S.~Zapperi, J.~Weiss, and J.~R. Grasso,
  ``Complexity in dislocation dynamics model,'' {\em Mater. Sci. Engr. A},
  vol.~309--310, pp.~324--327, 2001.

\bibitem{Miguel2002}
M.~C. Miguel, A.~Vespignani, M.~Zaiser, and S.~Zapperi, ``Dislocation jamming
  and {A}ndrade creep,'' {\em Phys. Rev. Letters}, vol.~89, p.~165501, 2002.

\bibitem{MiguelMZZ05}
M.~Miguel, P.~Moretti, M.~Zaiser, and S.~Zapperi, ``Statistical dynamics of
  dislocations in simple models of plastic deformation phase transitions and
  related phenomena,'' {\em Materials Science Engineering A}, vol.~400,
  pp.~191--198, 2005.

\bibitem{MiguelNature2001}
M.~C. Miguel, A.~Vespignani, S.~Zapperi, J.~Weiss, and J.-R. .~R. Grasso,
  ``Intermittent dislocation flow in viscoplastic deformation,'' {\em Nature},
  vol.~410, no.~6829, p.~667, 2001.

\bibitem{MiguZapp06}
M.~C. Miguel and S.~Zapperi, ``Fluctuations in plasticity at the microscale,''
  {\em Science}, vol.~312, pp.~1151--52, 2006.

\bibitem{MorettiMZZ04}
P.~Moretti, M.-C. Miguel, M.~Zaiser, and S.~Zapperi, ``Depinning transition of
  dislocation assemblies pileups and low-angle grain boundaries,'' {\em Phys.
  Rev. B}, vol.~69, no.~21, p.~214103, 2004.

\bibitem{nukala2003}
P.~K. V.~V. Nukala and S.~Simunovic, ``An efficient algorithm for simulating
  fracture using large fuse networks,'' {\em J. Phys. A Math. Gen.}, vol.~36,
  no.~45, pp.~11403--11412, 2003.

\bibitem{nukala2004percolation}
P.~Nukala, S.~{v{S}}imunovi{c}, and S.~Zapperi, ``Percolation and localization
  in the random fuse model,'' {\em Journal of Statistical Mechanics Theory and
  Experiment}, vol.~2004, p.~P08001, 2004.

\bibitem{nukala2005}
P.~K. V.~V. Nukala and S.~ifmmode check{S}else
  v{S}fi{}imunoviifmmode~acute{c}else {c}fi{}, ``Statistical physics models for
  nacre fracture simulation,'' {\em Phys. Rev. E}, vol.~72, p.~041919, 2005.

\bibitem{nukala2005statistical}
P.~Nukala, S.~Zapperi, and S.~{v{S}}imunovi{c}, ``Statistical properties of
  fracture in a random spring model,'' {\em Physical Review E}, vol.~71, no.~6,
  p.~66106, 2005.

\bibitem{nukala20101}
P.~K. V.~V. Nukala, P.~Barai, and R.~Sampath, ``An algorithm for simulating
  fracture of cohesivefrictional materials,'' {\em Journal of Statistical
  Mechanics Theory and Experiment}, vol.~2010, no.~11, p.~P11004, 2010.

\bibitem{NukalaZS06}
P.~K. V.~V. Nukala, S.~Zapperi, and S.~v{S}imunovi{c}, ``Crack surface
  roughness in three-dimensional random fuse networks,'' {\em Physical Review E
  Statistical, Nonlinear, and Soft Matter Physics}, vol.~74, no.~2, p.~026105,
  2006.

\bibitem{PapanikolaouBSDZS11}
S.~Papanikolaou, F.~Bohn, R.~L. Sommer, G.~Durin, S.~Zapperi, and J.~P. Sethna,
  ``Universality beyond power laws and the average avalanche shape,'' {\em
  Nature Physics}, vol.~7, pp.~316--320, 2011.

\bibitem{PhysRevE.61.R3283}
M.~Barthelemy, S.~V. Buldyrev, S.~Havlin, and H.~E. Stanley, ``Multifractal
  properties of the random resistor network,'' {\em Phys. Rev. E}, vol.~61,
  pp.~R3283--R3286, 2000.

\bibitem{PhysRevLett.105.155502}
C.~B. Picallo, J.~M. Lopez, S.~Zapperi, and M.~J. Alava, ``From brittle to
  ductile fracture in disordered materials,'' {\em Phys. Rev. Lett.}, vol.~105,
  p.~155502, 2010.

\bibitem{pradhan2005crossover}
S.~Pradhan, B.~Chakrabarti, and A.~Hansen, ``Crossover behavior in a mixed-mode
  fiber bundle model,'' {\em Physical Review E}, vol.~71, no.~3, p.~36149,
  2005.

\bibitem{pradhan2005crossoverI}
S.~Pradhan, A.~Hansen, and P.~Hemmer, ``Crossover behavior in burst avalanches
  signature of imminent failure,'' {\em Physical review letters}, vol.~95,
  no.~12, p.~125501, 2005.

\bibitem{pradhan2009}
S.~Pradhan, A.~Hansen, and B.~K. Chakrabarti, ``Failure processes in elastic
  fiber bundles,'' {\em Rev. Mod. Phy.}, 2009.

\bibitem{raisanen1998fracture}
V.~Raisanen, M.~Alava, and R.~Nieminen, ``Fracture of three-dimensional fuse
  networks with quenched disorder,'' {\em Physical Review B}, vol.~58, no.~21,
  pp.~14288--14295, 1998.

\bibitem{raisanen1998quasistatic}
V.~Raisanen, E.~Seppala, M.~Alava, and P.~Duxbury, ``Quasistatic cracks and
  minimal energy surfaces,'' {\em Physical review letters}, vol.~80, no.~2,
  pp.~329--332, 1998.

\bibitem{ramstad2004}
T.~Ramstad, J.~O.~H. Bakke, J.~Bjelland, T.~Stranden, and A.~Hansen,
  ``Correlation length exponent in the three-dimensional fuse network,'' {\em
  Phys. Rev. E}, vol.~70, p.~036123, 2004.

\bibitem{shekhawat2013fracture}
A.~Shekhawat, {\em FRACTURE IN DISORDERED BRITTLE MEDIA}.
\newblock PhD thesis, Cornell University, 2013.

\bibitem{ravi1994}
R.~Bhagavatula, K.~Chen, C.~Jayaprakash, and H.-J. Xu, ``Greens function method
  for random fuse network problems,'' {\em Phys. Rev. E}, vol.~49,
  pp.~5001--5006, 1994.

\bibitem{roux1988}
S.~Roux, A.~Hansen, H.~Herrmann, and E.~Guyon, ``Rupture of heterogeneous media
  in the limit of infinite disorder,'' {\em Journal of Statistical Physics},
  vol.~52, pp.~237--244, 1988.

\bibitem{Roux2011}
M.~Talamali, V.~Petaja, D.~Vandembroucq, and S.~Roux, ``Avalanches, precursors,
  and finite-size fluctuations in a mesoscopic model of amorphous plasticity,''
  {\em Phys. Rev. E}, vol.~84, p.~016115, 2011.

\bibitem{roux87}
J.-N. Roux and D.~Wilkinson, ``Resistance jumps in mercury injection in porous
  media,'' {\em Phys. Rev. A}, vol.~37, no.~10, pp.~3921--3926, 1988.

\bibitem{roux91}
S.~Roux, A.~Hansen, and E.~L. Hinrichsen, ``Multifracctality of conductance
  jumps in percolation,'' {\em Phys. Rev. B}, vol.~43, p.~3601, 1991.

\bibitem{shekhawat2011}
A.~Shekhawat, S.~Papanikolaou, S.~Zapperi, and J.~P. Sethna, ``Dielectric
  breakdown and avalanches at nonequilibrium metal-insulator transitions,''
  {\em Physical Review Letters}, vol.~107, no.~27, p.~276401, 2011.

\bibitem{shekhawat2013}
A.~Shekhawat, S.~Zapperi, and J.~P. Sethna, ``From damage percolation to crack
  nucleation through finite size criticality,'' {\em Phys. Rev. Lett.},
  vol.~110, p.~185505, 2013.

\bibitem{shekhawat2014}
Z.~Bertalan, A.~Shekhawat, J.~P. Sethna, and S.~Zapperi, ``Fracture strength
  stress concentration, extreme value statistics, and the fate of the weibull
  distribution,'' {\em Phys. Rev. Applied}, vol.~2, p.~034008, 2014.

\bibitem{sornette1992dynamics}
D.~Sornette and C.~Vanneste, ``Dynamics and memory effects in rupture of
  thermal fuse networks,'' {\em Physical review letters}, vol.~68, no.~5,
  pp.~612--615, 1992.

\bibitem{sornette2000}
A.~Johansen and D.~Sornette, ``Critical ruptures,'' {\em The European Physical
  Journal B - Condensed Matter and Complex Systems}, vol.~18, no.~1,
  pp.~163--181, 2000.

\bibitem{sornette2001}
S.~Gluzman and D.~Sornette, ``Self-consistent theory of rupture by progressive
  diffuse damage,'' {\em Phys. Rev. E}, vol.~63, p.~066129, 2001.

\bibitem{springerlink10.1007s1166101002925}
N.~Hansen, X.~Huang, and G.~Winther, ``Effect of grain boundaries and grain
  orientation on structure and properties,'' {\em Metallurgical and Materials
  Transactions A}, vol.~42, no.~3, pp.~613--625, 2011.

\bibitem{vanneste1992dynamical}
C.~Vanneste and D.~Sornette, ``The dynamical thermal fuse model,'' {\em Journal
  de Physique I}, vol.~2, no.~8, pp.~1621--1644, 1992.

\bibitem{WeissGrasso2001}
J.~Weiss, J.-R. Grasso, M.-C. Miguel, A.~Vespignani, and S.~Zapperi,
  ``Complexity in dislocation dynamics experiments,'' {\em Materials Science
  and Engineering A}, vol.~309310, no.~0, pp.~360 -- 364, 2001.

\bibitem{wu2004}
F.~Y. Wu, ``Theory of resistor networks the two-point resistance,'' {\em J.
  Phys. A Math. Gen.}, vol.~37, pp.~6653--6673, 2004.

\bibitem{zapperi05}
S.~Zapperi, P.~K. V.~V. Nukala, and S.~Simunovic, ``Crack roughness and
  avalanche precursors in the random fuse model,'' {\em Phys. Rev. E}, vol.~71,
  p.~026106, 2005.

\bibitem{zapperi05b}
S.~Zapperi, P.~K. V.~V. Nukala, and S.~Simunovic, ``Crack avalanches in the
  three-dimensional random fuse model,'' {\em Physica A}, vol.~357, p.~129,
  2005.

\bibitem{zapperi1997}
S.~Zapperi, P.~Ray, H.~E. Stanley, and A.~Vespignani, ``First-order transition
  in the breakdown of disordered media,'' {\em Phys. Rev. Lett.}, vol.~78,
  pp.~1408--1411, 1997.

\bibitem{zapperi1997plasticity}
S.~Zapperi, A.~Vespignani, and H.~Stanley, ``Plasticity and avalanche behaviour
  in microfracturing phenomena,'' {\em Nature}, vol.~388, no.~6643,
  pp.~658--659, 1997.

\bibitem{zapperi1999analysis}
S.~Zapperi, P.~Ray, H.~Stanley, and A.~Vespignani, ``Analysis of damage
  clusters in fracture processes,'' {\em Physica A Statistical Mechanics and
  its Applications}, vol.~270, no.~1-2, pp.~57--62, 1999.

\bibitem{zapperi1999avalanches}
S.~Zapperi, P.~Ray, H.~Stanley, and A.~Vespignani, ``Avalanches in breakdown
  and fracture processes,'' {\em Physical Review E}, vol.~59, no.~5,
  pp.~5049--5057, 1999.

\bibitem{zapperi2000planar}
S.~Zapperi, H.~Herrmann, and S.~Roux, ``Planar cracks in the fuse model,'' {\em
  The European Physical Journal B-Condensed Matter and Complex Systems},
  vol.~17, no.~1, pp.~131--136, 2000.

\end{thebibliography}

\documentclass[a4paper]{article}
\usepackage{geometry,graphicx,amsmath,amssymb,cite}
\title{Planar 2-D Cracks And Inclusions In Elastic Media}
\author{Raghu Singh Rathore\\ Jaipur Institute of Technology}
\begin{document}
\maketitle
\begin{abstract}
Electrical analogues of fracture, such as the fuse network model, are widely studied. However, the 
``analogy'' between the electrical problem and the elastic problem is rarely established explicitly. 
Further, the fuse network is a discrete approximation to the continuous problem of fracture. 
It is rarely, if ever, shown that the discrete approximation indeed approaches its continuum limit. 
We establish both of these correspondences directly.
\end{abstract}
\section{Introduction}
The fuse network is a widely studied model of brittle fracture~\cite{sornette1987, 0022-3727-42-21-214014, alava09, Alava2002, alava2006, alava2006morphology, alava2008fracture, alava2008role, Alava2010, Anup2015168, arcangelis85, bardley1994, bastiaansen97, beale1988, bergman1981, blackman76, bour2002statistical, BrownHCMLSC04, ChenPSZD11, shekhawat2012ProcI,colina1993model, Csikor2007, curtin1990brittle, davy1995localization, de1989scaling, shekhawat2012ProcII,de1989scalingI, delaplace1996progressive, duxbury1986, duxbury1987, duxbury1987failure, duxbury1994, duxbury1994failure, shekhawat2012ProcIII,duxbury1994size, hansen1991, hansen1991, hansen1991roughness, hansen2003, hansen2005, hansen2006, hansen2012, hemmer1992, hemmer1997, hidalgo2001bursts, shekhawat2014Imp, Hope20094593, HughChrzLiu98, HughesActaMater1997, HughesPRL2001, HughesScrMater2003, JohansenApJ2006, Kahng1988, korteoja1998, KuhlmannWilsdorfScrMetalMater1991, kun2000damage, LaPortaZS12, LaursonA2006, LuH2009, LuoZHL2012, MagniDZS09, MagniDZSnn, ManzatoSNASZ12, shekhawat2013ProcI,Miguel2001, Miguel2002, MiguelMZZ05, MiguelNature2001, MiguZapp06, MorettiMZZ04, nukala2003, nukala2004percolation, nukala2005, nukala2005statistical, nukala20101, NukalaZS06, PapanikolaouBSDZS11, PhysRevE.61.R3283, PhysRevLett.105.155502, pradhan2005crossover, pradhan2005crossoverI, pradhan2009, raisanen1998fracture, raisanen1998quasistatic, ramstad2004, shekhawat2013fracture,ravi1994, roux1988, Roux2011, roux87, roux91, shekhawat2011, shekhawat2013, shekhawat2014, sornette1992dynamics, sornette2000, sornette2001, springerlink10.1007s1166101002925, vanneste1992dynamical, WeissGrasso2001, wu2004, zapperi05, zapperi05b, zapperi1997, zapperi1997plasticity, zapperi1999analysis, zapperi1999avalanches, zapperi2000planar}. 
It is claimed that a fuse network is a discrete approximation to the mode III problem of fracture in a linear elastic isotropic continuum. However, this claim is rarely established with any 
rigor. The fuse network has also been used as a discrete approximation to a conduction problem, where an inclusion with conductivity mismatch is embedded in a uniform medium. 
Here we establish both of these correspondences in a rigorous manner and show the behavior in some singular limits, such as those of a long sharp crack, and a long sharp inclusion.
\par
The early work on fuse networks concentrated on the strength distribution of diluted fuse networks~\cite{duxbury1986,duxbury1987,duxbury1987failure,duxbury1994,duxbury1994failure}, 
and was led by Duxbury and co-workers. The diluted fuse networks were weakly disordered; a small fraction of their fuses were removed in the beginning. 
Due to the small fraction, most of the removed fuses were isolated from each other. However, random statistical fluctuations lead to configurations
where a contiguous row to fuses is removed. If the dilution probability is $p$, then the probability of having a contiguous row of $n$ 
removed fuses is $p^n$. Since there are $L^2$ lattice sites in a 2D square network of size $L$, therefore, the expected number 
of cracks of size $n$ in the network scales as $L^2 p^n$. Thus, one can find the expected length of the longest crack by 
setting $L^2 p^n \sim 1$ or $n \sim -2\log L/\log p$. Since the `stress concentration' at the tip of the crack scales as $\sqrt{n}$, one can work out the 
expected strength of the network. A vast literature is devoted towards understanding the strength of such networks. 
Recently, the asymptotic behavior of the strength distribution has been studied in context of extreme value statistics~\cite{ManzatoSNASZ12,shekhawat2014Imp}.
\par
The more modern literature deals with the problem of distributed strengths, as in where the strength of each individual fuse is a random variable 
taken from a certain distribution. The effort on this front has been led by Zapperi and 
co-workers~\cite{zapperi05,zapperi05b,zapperi1997,zapperi1997plasticity,zapperi1999analysis,zapperi1997,shekhawat2013,shekhawat2014}, and
Hansen and co-workers~\cite{hansen2012,hansen1991,hansen1991roughness,hansen2003,hansen2005, roux87,roux1988,roux1988,arcangelis85}. 
The disorder in this problem can be made strong by choosing a distribution with a large variance, or a large spread of the values by having a broad spectrum (typically, $P(X_i<x) = x^\beta$,
where $X_i$ is the strength of the $i^{th}$ fuse, and the distribution becomes broad in the limit of $\beta \to 0$). 
This problem exhibits elements of scale-invariance, entangled with elements of first order nucleation~\cite{shekhawat2013}.
\par
In the study of all of these problems it is necessary to understand the behavior of an isolated crack in the fuse network.
While this problem has been studied and solved before, its hard to find the solution, and reproducing it is not a trivial exercise. 
Here we reproduce the solution to the problem of an elliptical inclusion of a mismatched conductance in an otherwise uniform infinite 2D medium.
A sharp crack (or sharp needle) is obtained by taking the limit of a thin ellipse, and setting the conductance mismatch appropriately. 
Further, the lattice limit is approached by integrating the continuum solution over one lattice spacing.
\section{Mathematical Preliminaries}
\subsection{Elliptical Coordinate System}
The elliptical coordinate system is related to the Cartesian coordinate system by the following transformations
\begin{align}
x & = c \cosh \xi \cos \eta, \\
y & = c \sinh \xi \sin \eta,
\label{eq:EllipDefn}
\end{align}
where $c > 0$ is a parameter. Setting
\begin{equation}
c = \left( a^2 - b^2 \right)^{1/2},
\end{equation}
yields the ellipse $x^2/a^2 + y^2/b^2 = 1$ as the curve traced by $(\xi_0, \eta)$, where
\begin{align}
a & = c \cosh \xi_0,\\
b & = c \sinh \xi_0.
\end{align}
Defining the aspect ratio of the ellipse as 
\begin{equation}
n \equiv \frac{a}{b},
\end{equation}
one gets
\begin{align}
e^{\xi_0} & =  \left( \frac{n+1}{n-1} \right )^{1/2},\label{eq:EXI}\\
\mathrm{or}\ \xi_0 & = \frac{1}{2} \ln \left( \frac{n+1}{n-1} \right ) \label{eq:XI}.
\end{align}

\subsection{Jacobian Matrix}
The Jacobian matrix of the transformation defined by Eq.~\ref{eq:EllipDefn} is given by
\begin{align}
\mathbf{J} & = \left( \begin{array}{cc}
			\frac{\partial x}{\partial \xi}	& \frac{\partial y}{\partial \xi}\\
			\frac{\partial x}{\partial \eta}	& \frac{\partial y}{\partial \eta}
			\end{array} \right) \\
	 & = \left( \begin{array}{cc}
			c\sinh \xi \cos \eta 	& c \cosh \xi \sin\eta	\\
			-c\cosh \xi \sin \eta 	& c \sinh \xi \cos\eta
			\end{array} \right), 
\end{align}
and the determinant of the Jacobian is given by
\begin{equation}
\det( \mathbf{J}) = c^2(\sinh^2 \xi + \sin^2 \eta) = c^2( \cosh^2 \xi - \cos ^2 \eta)
\end{equation}

\subsection{Behavior Far From Origin}
For $r \to \infty$ the following approximations can be obtained
\begin{align}
	e^\xi & \to \frac{2r^2}{c^2}\ \mathrm{or}\ \xi \to \ln \left( \frac{2r^2}{c^2} \right), \\
	\eta & \to \theta,
\end{align}
where $x = r\cos \theta,\ y = r \sin \theta$.
\subsection{Behavior Near Tip Of Thin Ellipses}
For thin ellipses $n \gg 1$, therefore, Eq.~\ref{eq:XI} yields
\begin{equation}
\xi_0 \approx \frac{1}{n} \ll 1.
\end{equation}
\subsubsection{Behavior Along Major Axis}
Near the tip of a thin ellipse $\xi \ll 1$, and, $\eta = 0$ along the major axis. 
Taking $x = a + \delta b,\ y = 0$, we get
\begin{equation}
a+b\delta = c\cosh \xi \approx c\left( 1 + \frac{\xi^2}{2} \right).
\end{equation}
Inverting the above yields
\begin{equation}
\xi^2 \approx 2\left( \frac{a+b\delta}{c} - 1\right) \approx 2\left( \frac{\delta}{n} \right),
\end{equation}
where the last approximation is valid since for $n \gg 1$, $c \approx a$. Thus, we have
\begin{equation}
\xi \approx \left(\frac{2\delta}{n}\right)^{1/2}
\label{eq:LatticeStepHor}
\end{equation}
\subsubsection{Behavior Parallel To Minor Axis}
Taking $x = a$, $y = \delta b$, we get
\begin{align}
a & = c\cosh \xi \cos \eta \\
\delta b & = c\sinh \xi \sin \eta.
\end{align} 
Solving the above in the limit of $n \gg 1$, gives
\begin{equation}
\xi \approx \left(\frac{\delta}{n}\right)^{1/2} \approx \eta
\label{eq:LatticeStepVer}
\end{equation}

\subsection{Gradient}
The general expression for the gradient is 
\begin{equation}
\nabla f = \sum_i \frac{1}{h_i}\frac{f}{q_i} {q}_i,
\end{equation}
where $q_i$ are the coordinates, $\hat{q}_i$ are the unit vectors, and $h_i$ are defined as follows
\begin{equation}
h^2_i = \sum_k \left( \frac{\partial X_k}{\partial q_i} \right)^2,
\end{equation}
where $\vec{r} = X_k \hat{e}_k$, $\hat{e}_k$ being the Cartesian unit vectors. For the elliptical coordinates we have
\begin{equation}
\vec{r} = c\cosh \xi \cos \eta \hat{e}_1 + c \sinh \xi \sin \eta \hat{e}_2.
\end{equation}
Thus, we get
\begin{equation}
\begin{split}
\nabla f = & \phantom{+} \frac{1}{h^2}\left( c\sinh \xi \cos \eta \frac{\partial f}{\partial \xi} - c\cosh \xi \sin \eta \frac{\partial f}{\partial \eta} \right)\hat{e}_1 \\ 
	 & + \frac{1}{h^2}\left( c\cosh \xi \sin \eta \frac{\partial f}{\partial \xi} + c\sinh \xi \cos \eta \frac{\partial f}{\partial \eta} \right)\hat{e}_2,
\end{split}
\label{eq:Grad}
\end{equation}
where 
\begin{equation}
h^2 = h^2_1 = h_2^2 = c^2(\cosh^2 \xi - \cos^2\eta). 
\end{equation}
\section{General Solution For Elastic Elliptical Inclusion In Elastic Media}
We consider a medium of conductance $\epsilon_1$, with an elliptical inclusion of conductance $\epsilon_2$. The inclusion is centered at the origin and 
has a major axis of length $a$, and a minor axis of length $b$. The major axis is oriented along the $x$-axis. The applied far-field voltage is $V_x x + V_yy$.
We solve the problem in an elliptical coordinate system defined by
\begin{align}
x & = c\cosh \xi \cos \eta, \\
y & = c\sinh \xi \sin \eta,
\end{align}
where $ c = (a^2 - b^2)^{1/2}$. The governing equation to be solved is the Laplace equation
\begin{equation}
\nabla^2 V^{in} = \nabla^2 V^{out} = 0,
\end{equation}
where $V^{in}(\xi,\eta)$ is the voltage field inside the inclusion, and $V^{out}(\xi,\eta)$ is the voltage field outside the inclusion.
The boundary conditions to be imposed are
\begin{align}
\lim_{\xi \to \infty} V^{out}(\xi,\eta) &= V_x x + V_y y,\label{eq:BC1} \\
\left. \epsilon_1 \frac{\partial V^{out}}{\partial \xi}\right|_{\xi = \xi_0} & = \left. \epsilon_2 \frac{\partial V^{in}}{\partial \xi}\right|_{\xi = \xi_0},\label{eq:BC2}\\
V^{out}(\xi_0,\eta) & = V^{in}(\xi_0,\eta)\label{eq:BC3}.
\end{align}
The Laplace equation is invariant under a coordinate transformation from the Cartesian to the elliptical coordinates. One can use the method of separation of 
variables to obtain the following solution
\begin{align}
V^{in}(\xi,\eta) & = \frac{n+1}{R+n}V_x x + \frac{n+1}{Rn+1}V_y y,\label{eq:Vin}\\
V^{out}(\xi,\eta) & = V_x x + V_y y + a V_x \left( \frac{1-R}{R+n} \right) e^{(\xi_0 - \xi)}\cos\eta + bV_y \left( \frac{1-R}{R + 1/n} \right) e^{(\xi_0 - \xi)}\sin\eta\label{eq:Vout}.
\end{align}
Instead of deriving the above solution, we simply prove that it is a valid solution to the posed problem, and then appeal to the uniqueness of solution 
of Laplace equations. It is easy to see that the boundary condition given by Eq.~\ref{eq:BC1} is satisfied by the proposed solution. The boundary condition 
given by Eq.~\ref{eq:BC2} is satisfied since
\begin{align*}
\epsilon_2 \frac{\partial V^{in}}{\partial \xi} & = \epsilon_2 \frac{n+1}{R+n}V_x c\sinh \xi \cos \eta + \epsilon_2 \frac{n+1}{Rn+1}V_y c\cosh\xi \sin\eta\\
\Rightarrow \epsilon_2 \left. \frac{\partial V^{in}}{\partial \xi}\right|_{\xi = \xi_0} & = \epsilon_2 \frac{n+1}{R+n}V_x b\cos \eta + \epsilon_2\frac{n+1}{Rn+1}V_y a \sin\eta
\end{align*}
and
\begin{align*}
\epsilon_1 \frac{\partial V^{out}}{\partial \xi} & = \epsilon_1 V_x c\sinh \xi \cos \eta + \epsilon_1  V_y c\cosh\xi \sin\eta\\
 & \phantom{=} - \epsilon_1 a V_x \left( \frac{1-R}{R+n} \right) e^{(\xi_0 - \xi)}\cos\eta - \epsilon_1 bV_y \left( \frac{1-R}{R + 1/n} \right) e^{(\xi_0 - \xi)}\sin\eta \\
\Rightarrow \epsilon_1 \left. \frac{\partial V^{out}}{\partial \xi}\right|_{\xi=\xi_0} & = \epsilon_1 V_x b \cos \eta + \epsilon_1  V_y a \sin\eta
 -\epsilon_1 a V_x \left( \frac{1-R}{R+n} \right) \cos\eta - \epsilon_1 bV_y \left( \frac{1-R}{R + 1/n} \right) \sin\eta \\
	& = \epsilon_1 \left( b - a \frac{1-R}{R+n} \right )V_x\cos\eta + \epsilon_1  \left(a -b \frac{1-R}{R + 1/n}\right)V_y\sin\eta\\
	& = \epsilon_1R \frac{a+b}{R+n} V_x\cos\eta + \epsilon_1 R \frac{a+b}{R + 1/n}V_y\sin\eta \\
	& = \epsilon_2 \frac{n+1}{R+n}V_x b\cos \eta + \epsilon_2\frac{n+1}{Rn+1}V_y a \sin\eta \\
	& = \epsilon_2 \left. \frac{\partial V^{in}}{\partial \xi}\right|_{\xi = \xi_0} 
\end{align*}
Finally, the boundary condition given by Eq.~\ref{eq:BC3} is satisfied since
\begin{align*}
V^{out}(\xi_0,\eta) & = V_x a\cos\eta + V_y b\sin\eta + a V_x \left( \frac{1-R}{R+n} \right) \cos\eta + bV_y \left( \frac{1-R}{R + 1/n} \right) \sin\eta\\
	& = \frac{n+1}{R+n}V_x a\cos\eta + \frac{1+1/n}{R+1/n}V_y b\sin \eta\\
	& = V^{in}(\xi_0,\eta) 
\end{align*}
\section{Fracture}
The voltage field for an elliptical crack in a conductive medium can be found by setting $R = 0$ in Eq.~\ref{eq:Vout}. It is also customary to set
$V_x = 0$, so that the applied field is perpendicular to the crack. Thus, for cracks we get
\begin{equation}
V^{cr}(\xi,\eta)  = V_y y + bV_y n e^{(\xi_0 - \xi)}\sin\eta = V_y y + bV_y n\left(\frac{n+1}{n-1}\right)^{1/2} e^{-\xi}\sin\eta \label{eq:Vcr},
\end{equation}
where the last equality follows from Eq.~\ref{eq:EXI}. By using Eq.~\ref{eq:Grad} the gradient of the voltage field can be found to be
\begin{align}
\nabla V^{cr}(\xi,\eta) = & \phantom{.....} \frac{1}{h^2}\left( c\sinh \xi \cos \eta \frac{\partial V^{cr}}{\partial \xi} - c\cosh \xi \sin \eta \frac{\partial V^{cr}}{\partial \eta} \right)\hat{e}_1 \\ 
	 & + \frac{1}{h^2}\left( c\cosh \xi \sin \eta \frac{\partial V^{cr}}{\partial \xi} + c\sinh \xi \cos \eta \frac{\partial V^{cr}}{\partial \eta} \right)\hat{e}_2\\
	 =& -V_y\frac{cbn e^{\xi_0}}{2h^2}\sin( 2\eta) \hat{e}_1 + V_y \hat{e}_2 + V_y\frac{cbne^{\xi_0}}{2h^2}(\cos 2\eta - e^{-2\xi})\hat{e}_2 \label{eq:FracGrad}
\end{align}
\subsection{Lattice Limit}
The current density in a continuous media is given by $\epsilon \nabla V$, where $\epsilon$ is the conductivity of the media. However, the fuse networks are a 
lattice approximation to the continuous media, and thus to find the current in a fuse adjacent to a crack, one has to take an adequate lattice limit 
of Eq.~\ref{eq:FracGrad}. A straight `crack' of length $n\beta$ in a fuse network comprises of a row of $n$ burned fuses in a straight line, where $\beta$ is 
one lattice constant. This crack can be approximated by an ellipse of length $n\beta$ and width $\beta$, thus, the aspect ratio of the ellipse is 
equal to $n$, the semi-major axis, $a$, is equal to $n\beta/2$ and the semi-minor axis, $b$, is equal to $\beta/2$. 
To find the current in the bond adjacent to the crack one has to integrate the current density, $\epsilon\nabla V$, over one lattice constant.
We work in the limit of long, thin cracks, i.e.~$n \gg 1$. Taking $\eta = 0$ ahead of an elliptical crack, Eq.~\ref{eq:FracGrad} yields
\begin{align*}
\nabla V^{cr}(\xi,\eta = 0) =& V_y \hat{e}_2 + V_y\frac{bne^{\xi_0}}{2c(\cosh^2\xi - 1)}(1 - e^{-2\xi})\hat{e}_2,\\ 
		\approx & V_y \hat{e}_2 + V_y\frac{bn}{c\xi} \hat{e}_2,\\
		\approx & V_y \hat{e}_2 + V_y\left(\frac{n}{2\delta}\right)^{1/2} \hat{e}_2,
\end{align*}
where, we have used $\xi\approx \sqrt{2\delta/n} \ll 1$ ahead of a thin ellipse, with $x = a+\delta b,\ y = 0$, $0 < \delta < 2$. Thus, the current
at the tip of crack of length $n$ is given by
\begin{equation}
I^{tip}_n  \approx \epsilon \int_0^2 (\nabla V^{cr}\cdot \hat{e}_2)b d\delta \approx \epsilon V_y\beta(1 + \sqrt{n}).
\label{eq:FracTip}			
\end{equation}
The above result shows that the current enhancement at the tip of a crack of length $n$ in a fuse network is proportional to $\sqrt{n}$.
\section{Dielectric Breakdown}
For the case of dielectric breakdown, the electric field is taken to be aligned with major axis of the elliptical inclusion. Thus, we set
$V_y = 0$ in Eq.~\ref{eq:Vout} to get
\begin{equation}
V^{db}(\xi,\eta) = V_x x + a V_x \left( \frac{1-R}{R+n} \right) e^{(\xi_0 - \xi)}\cos\eta 
\label{eq:db}
\end{equation}
One also need to take the limit of $R\to \infty$ for a highly conductive inclusion, but we leave $R$ finite for now.
As before, taking the gradient of the above gives
\begin{equation}
\nabla V^{db}(\xi,\eta) = V_x\hat{e}_1 + V_x\frac{ac e^{\xi_0}}{2h^2}\left( \frac{R-1}{R+n}\right)(\cos 2\eta - e^{-2\xi})\hat{e}_1
+ V_x\frac{ac e^{\xi_0}}{2h^2}\left( \frac{R-1}{R+n}\right) \cos 2\eta\hat{e}_2
\label{eq:ElecGrad}
\end{equation}
\subsection{Lattice Limit}
As before, we take the lattice limit by considering an ellipse of length $n\beta$ and width $\beta$, where $\beta$ is the lattice constant.
We find the tip current by integrating the current density perpendicular to the contour $x = a,\ y = \delta b$, $-1 < \delta < 1$. By using 
Eq.~\ref{eq:LatticeStepVer} we get, along the contour,
\begin{align*}
\nabla V^{db}\cdot \hat{e}_1 & \approx V_x + \frac{V_x}{2}\left(\frac{n}{\delta}\right)^{1/2}\frac{R-1}{R+n}.
\end{align*}
Thus, the current at the tip of a conducting inclusion is given by
\begin{equation}
I^{tip}_n \approx 2\epsilon \int_0^1 (\nabla V^{db} \cdot \hat{e}_1)bd\delta \approx \epsilon V_x \beta(1 + \frac{R-1}{R+n}\sqrt n)
\label{eq:BoltTip}
\end{equation}
Thus, for $R\to \infty$, the current enhancement at the tip of a conducting bolt is proportional to $\sqrt n$; however, for any finite 
$R$, the current enhancement dies as $(R-1)/\sqrt n$.
\bibliographystyle{ieeetr}
\bibliography{bibfile}
\end{document}